# On Reducing the Execution Latency of Superconducting Quantum Processors via Quantum Program Scheduling

### v1

@@ Abstract.
Quantum computing has gained considerable attention, especially after the arrival of the Noisy Intermediate-Scale Quantum (NISQ) era. Quantum processors and cloud services have been made world-wide increasingly available. Unfortunately, programs on existing quantum processors are often executed in series, and the workload could be heavy to the processor. Typically, one has to wait for hours or even longer to obtain the result of a single quantum program on public quantum cloud due to long queue time. In fact, as the scale grows, the qubit utilization rate of the serial execution mode will further diminish, causing the waste of quantum resources. In this paper, to our best knowledge for the first time, the Quantum Program Scheduling Problem (QPSP) is formulated and introduced to improve the utility efficiency of quantum resources. Specifically, a quantum program scheduling method concerning the circuit width, number of measurement shots, and submission time of quantum programs is proposed to reduce the execution latency. We conduct extensive experiments on a simulated Qiskit noise model, as well as on the Xiaohong (from QuantumCTek) superconducting quantum processor. Numerical results show the effectiveness in both QPU time and turnaround time.
copyright: none
@@ 1. Introduction
More and more players like IBM have provided the public access to their quantum computers. This facilitates the validation of quantum algorithms on NISQ devices over the Internet. For example, we have free access to the 7-qubit IBM Perth (Computing, 2019). However, running quantum programs on current quantum computers is non-trivial due to the noise and sparse connectivity of physical qubits. On a NISQ device, the physical qubits are not fully connected. The deployment of two-qubit gates is restricted to pairs of connected qubits. Hence, when mapping logical qubits to their physical counterparts, certain two-qubit gates may be positioned on physically disconnected qubits, rendering them inexecutable. Conventionally, SWAP gates are inserted to change the qubit mapping so that every two-qubit gate can be physically executed.
A more awkward obstacle hindering people from using quantum computers is the unbearably long queue time. Though there exist some quantum cloud services, the growing need for quantum hardware outpaces the open access to quantum hardware. We submit many jobs to IBM Perth. The average number of pending jobs is about 2,540, and the average queue time is about 6.7 hours. The latency of circuit execution is unacceptable, especially when we run Variational Quantum Algorithms (VQAs) (Cerezo et al., 2021), in which plenty of circuits are executed in a single episode to update the parameters. The main reason for this latency is that the submitted quantum programs are executed in series. In other words, only one program is executed on the quantum processor in each execution. Besides, the qubit utilization rate is low. Since entangling a large number of qubits on NISQ devices is challenging due to the noise (Cao et al., 2023), most circuits remain small in width to ensure high fidelity. With the increasing number of physical qubits on QPUs and decreasing error rate, we may execute multiple programs in parallel in each execution, i.e. multi-programming, at a negligible cost of fidelity to reduce the latency. As a result, more people can access quantum resources to facilitate QC.
@@
However, multi-programming on quantum processors is a complicated task. The execution order of programs will affect the performance of multi-programming. Different from classical process scheduling, we need to consider fidelity apart from time metrics. The QPU should be partitioned in a fair manner to reduce fidelity drop. Unfortunately, fidelity and time metrics often conflict with each other. In this paper, we introduce the Quantum Program Scheduling Problem, which has great practical value in the NISQ era. A novel scheduling method is proposed to tackle this problem. With our priority score and noise-aware initial mapping, our method surpasses baselines in time metrics, and guarantees the fairness and fidelity. The contributions of this paper are:
1) We formulate the Quantum Program Scheduling Problem of reducing the latency of superconducting quantum processors.
2) We propose a novel scheduling method to balance different metrics. Also, three greedy baselines are provided for comparison.
3) Experiments on both the noise model and quantum computer show that our approach significantly reduces the QPU time and turnaround time at a low cost of fidelity.
@@ 2. Preliminaries and Related Works
Quantum Computing. The basic unit in QC is a qubit, which is in superposition of basis states |$\ket{0}$ and |$\ket{1}$: |$\ket{\psi}=a\ket{0}+b\ket{1}$. Likewise, a quantum system with $n$ qubits is in superposition of $2^{n}$ basis states. The evolution of quantum states can be used to generate solutions to specific problems, perhaps in a much faster manner than classical methods. We refer readers to (Nielsen and Chuang, 2010) for detailed backgrounds. Similar to classical circuits, quantum circuits are employed to implement quantum computation of quantum states. Each quantum circuit consists of quantum gates like X gates, RZ gates, CNOT gates, etc. To obtain the result, we have to repeat executing the circuits many times (shots), because quantum measurements will cause the collapse of a superposition state to a basis state. In this paper, we denote the quantum circuits used to solve specific problems as quantum programs.
@@
Quantum Processors. The core of a quantum computer is the quantum processor, aka QPU, which serves to execute quantum circuits. We focus on superconducting quantum processors in this paper. The major properties of a QPU are its basis gates, coupling graph and noise condition. Basis gates are the quantum gates supported on the QPU. All the gates in a quantum program must be converted to basis gates before execution during compiling. As shown in Fig. 1, the coupling graph restricts the connectivity of qubits. Two-qubit gates can only be deployed on connected qubits. Besides, the noise of QPUs in the NISQ era results in gate errors, readout errors, and decoherence, which will corrupt the quantum state and reduce the fidelity. These errors change over time, so they must be calibrated regularly. Nowadays, many quantum processors are open to public through quantum cloud services. Our submitted quantum programs will queue up to be executed. In this paper, we conduct experiments on the Qiskit (Aleksandrowicz et al., 2019) noise model of 16-qubit IBM Guadalupe (Fig. 1b), and 66-qubit Xiaohong [^1] As used in our experiments, Xiaohong is a 66-qubit superconducting quantum processor, which can be accessed via public cloud at https://quantumctek-cloud.com/. The used QCIS instruction set can be easily converted from or to the widely used QASM. (Fig. 1a) quantum processor from the QuantumCTek Quantum Cloud Platform (Platform, 2022).
@@
Qubit Mapping. When logical qubits of a quantum program are mapped to physical qubits on a QPU, the original two-qubit gates may violate the connectivity constraints as shown in Fig. 1(b). A traditional way to solve this problem is to insert SWAP gates. A SWAP gate is implemented by three CNOT gates (Fig. 1c), incurring extra noise. Hence, the number of them is expected to be minimized. Siraichi et al. formally introduce the aforementioned qubit allocation (mapping) problem (Siraichi et al., 2018), which is proved to be NP-complete. Li et al. propose a bidirectional heuristic search (SABRE) to tackle this problem. When inserting a SWAP gate, they consider its impact on two-qubit gates in both the front layer and extended set, significantly reducing the SWAP overhead (Li et al., 2019). Niu et al. take the error rate and execution time of CNOT gates into consideration, and provide BRIDGE gates as an alternative to SWAP gates (Niu et al., 2020). The BRIDGE gate (Fig. 1d) is composed of four CNOT gates, but its effect equals a single CNOT gate. Niu et al. further ameliorate the mapping method by involving the cost of inserted SWAP gates and BRIDGE gates themselves (Niu and Todri-Sanial, 2023).
@@
Quantum Multi-Programming. Quantum multi-programming means running multiple quantum programs simultaneously on a QPU. Das et al. propose the concept of multi-programming NISQ devices to improve throughput. They allocate less noisy physical qubits to logical qubits with higher utility. Qucloud (Liu and Dou, 2021) leverages FN community detection algorithm (Newman, 2004) to partition QPUs, and designs an EPST score to estimate the fidelity of allocation. Resch et al. run multiple QAOA (Farhi et al., 2014) circuits in parallel to accelerate the training process (Resch et al., 2021). Niu et al. reorder the quantum programs according to their CNOT density, and partition QPUs based on the connectivity and error rates of physical qubits (Niu and Todri-Sanial, 2023). These existing methods either neglect the impact of execution order or just focus on multi-programming in single execution.
@@ 3. Methodology
In this section, we formally introduce the Quantum Program Scheduling Problem (QPSP) to excavate the importance of the execution order when multi-programming massive quantum programs in the queue of quantum cloud services. Also, a scheduling method and three greedy baselines are proposed to tackle this problem.
3.1. The Quantum Program Scheduling Problem
@@ 3.1.1. Definition
Suppose the current job queue $\mathcal{Q}$ is comprised of $K$ quantum jobs to be executed, i.e. $\mathcal{Q}=\{\mathcal{J}_{1},\mathcal{J}_{1},\cdots,\mathcal{J}_{K}\}$. Each job $\mathcal{J}_{i}$ can be represented as a tuple $(p_{i},s_{i},t_{i})$, where $p_{i}$, $s_{i}$, and $t_{i}$ denote the quantum program, number of measurement shots, and submission time, respectively. Then, new jobs $\mathcal{J}_{K+1},\mathcal{J}_{K+2},\cdots$ will be submitted at time $t_{K+1},t_{K+2},\cdots$. For a quantum computer, besides the execution time $t_{e}$ of circuits on the QPU, other procedures like circuit verification, generation of control signals, and communication will cost extra time $t_{m}$ between execution.
Given the coupling graph $\mathcal{G}$, noise calibration data $\mathcal{N}$, and basis gate set $\mathcal{B}$ of the QPU, we need to execute all the jobs submitted during a time period on the QPU. The objective of QPSP is to minimize the execution latency of jobs, and maintain high fidelity.
@@ 3.1.2. Metrics
The performance assessment of QPSP is divided into two parts: time and fidelity. Also, fairness should be considered.
Time. For users, they mainly care about the time cost from submission to completion of their quantum job, which we name turnaround time. For suppliers, they may put emphasis on the QPU time of their quantum processors, i.e. circuit running time on QPU. Besides, Trial Reduction Factor (TRF) (Das et al., 2019), which is the ratio of the number of trials when programs run in series to that when programs run in parallel, is included as a time metric in our paper.
Fidelity. The real fidelity of a quantum state is hard to obtain on a quantum computer. By convention, we use the Probability of Successful Trial (PST) (Tannu and Qureshi, 2019), which is defined as the percentage of trials producing the correct result, as our fidelity metric.
Figure 3. Overview of our method for scheduling.
@@ 3.2. Proposed Method
Our proposed method is composed of three parts: queue rearranging, qubit partitioning, and qubit mapping. Fig. 3 shows the overview of our method. In each iteration, we sort the current updated queue by our priority score $S_{p}$. Then, we evaluate quantum programs in the sorted queue one by one, selecting and mapping those programs that can find a partition on the remaining coupling graph, until the number of used physical qubits exceed the limit, i.e. $\Sigma_{i=1}^{k}n_{i}\leq\eta\cdot N$, where $n_{i}$ is the number of qubits (i.e. width) of the $i$-th selected programs, and $k$ is the number of selected programs. $N$ denotes the number of physical qubits, and $\eta\in(0,1]$ is the allowed maximum usage of physical qubits, which influences the average fidelity because higher usage will incur more noise. Then, the mapped programs will be executed in parallel on the QPU. The number of shots is set as the maximum shot number among the mapped programs to ensure all the shot requirements are satisfied, because those programs whose shot requirements are unsatisfied will lead to extra execution overhead in following iterations. The mapped programs are executed in an As Late As Possible (ALAP) manner so that programs with different depth can be completed at the same time to avoid decoherence.
@@ 3.2.1. Queue Rearranging
In each iteration, the program queue will be updated due to new submitted programs and executed programs. Akin to the importance of process scheduling for CPUs, the execution order of quantum programs also counts in QPSP. Therefore, we sort quantum programs in the updated queue in ascending order of their priority score. Three properties of a quantum job $\mathcal{J}_{i}$ are considered for its priority score $S_{p}^{(i)}$: the number of qubits $n_{i}$, number of shots $s_{i}$, and submission time $t_{i}$. The priority score is defined as the linear combination:
(1) $S_{p}^{(i)}=\alpha\cdot S_{n}^{(i)}+\beta\cdot S_{s}^{(i)}+\gamma\cdot S_{t}^{(i)},$
where $\alpha$, $\beta$, and $\gamma$ represent the width weight, shot weight, and time weight. $S_{n}^{(i)}$, $S_{s}^{(i)}$, and $S_{t}^{(i)}$ are the Min-max Normalization results of $n_{i}$, $s_{i}$, and $t_{i}$. For example, $S_{n}^{(i)}$ can be calculated as:
(2) $S_{n}^{(i)}=\frac{n_{i}-n_{min}}{n_{max}-n_{min}}.$

### v2

# On Reducing the Execution Latency of Superconducting Quantum Processors via Quantum Job Scheduling

@@ Abstract.
Quantum computing has gained considerable attention, especially after the arrival of the Noisy Intermediate-Scale Quantum (NISQ) era. Quantum processors and cloud services have been made world-wide increasingly available. Unfortunately, jobs on existing quantum processors are often executed in series, and the workload could be heavy to the processor. Typically, one has to wait for hours or even longer to obtain the result of a single quantum job on public quantum cloud due to long queue time. In fact, as the scale grows, the qubit utilization rate of the serial execution mode will further diminish, causing the waste of quantum resources. In this paper, to our best knowledge for the first time, the Quantum Job Scheduling Problem (QJSP) is formulated and introduced, and we accordingly aim to improve the utility efficiency of quantum resources. Specifically, a noise-aware quantum job scheduler (NAQJS) concerning the circuit width, number of measurement shots, and submission time of quantum jobs is proposed to reduce the execution latency. We conduct extensive experiments on a simulated Qiskit noise model, as well as on the Xiaohong (from QuantumCTek) superconducting quantum processor. Numerical results show the effectiveness in both the QPU time and turnaround time.
Quantum Computing, Quantum Job Scheduling, Quantum Cloud
journalyear: 2024copyright: acmlicensedconference: IEEE/ACM International Conference on Computer-Aided Design; October 27–31, 2024; New York, NY, USAbooktitle: IEEE/ACM International Conference on Computer-Aided Design (ICCAD ’24), October 27–31, 2024, New York, NY, USAdoi: 10.1145/3676536.3676678isbn: 979-8-4007-1077-3/24/10ccs: Hardware Quantum computation
@@ 1. Introduction
More and more players like IBM have provided the public access to their quantum computers. This facilitates the validation of quantum algorithms on NISQ devices over the Internet. For example, we have free access to the 7-qubit IBM Perth (Computing, 2019). However, running quantum circuits on current quantum computers is non-trivial due to the noise and sparse connectivity of physical qubits. On a NISQ device, the physical qubits are not fully connected. The deployment of two-qubit gates is restricted to pairs of connected qubits. Hence, when mapping logical qubits to their physical counterparts, certain two-qubit gates may be positioned on physically disconnected qubits, rendering them inexecutable. Conventionally, SWAP gates are inserted to change the qubit mapping so that every two-qubit gate can be physically executed. Since SWAP gates result in extra noise, the number of them is expected to be minimized.
A more awkward obstacle hindering people from using quantum computers is the unbearably long queue time. Though there exist some quantum cloud services, the growing need for quantum hardware outpaces the open access to quantum hardware. To verify this, we submit 20 jobs to IBM Perth within a week. According to the panel, the average number of pending jobs when submitting is about 2,540, and the average queue time before execution is about 6.7 hours. The latency of circuit execution is unacceptable, especially when we run Variational Quantum Algorithms (VQAs) (Cerezo et al., 2021), in which plenty of circuits are executed in a single episode to update the parameters. The main reason for this latency is that the submitted quantum jobs are executed in series. Thus, only one job is executed on the quantum processor in each execution. Besides, entangling a large number of qubits on NISQ devices is challenging due to the noise (Cao et al., 2023), so most circuits remain small in width to ensure high fidelity. Hence, the qubit utilization rate is low. With the increasing number of physical qubits on QPUs and decreasing error rate, we may execute multiple jobs in parallel in each execution, i.e. quantum multi-programming (QMP), at a negligible cost of fidelity to reduce the latency. Such parallel running mode can be extended to applications like quantum architecture search (Wu et al., 2023; Lu et al., 2023), quantum inner product (Xiong et al., 2024), and network alignment (Ye et al., 2023). As a result, more people can access quantum resources to facilitate QC.
@@
However, QMP on quantum processors is a complicated task. The execution order of circuits will affect its performance. Different from classical process scheduling, we need to consider fidelity apart from time metrics. The QPU should be partitioned in a fair manner to reduce fidelity drop. Unfortunately, fidelity and time metrics often conflict with each other. In this paper, we introduce the Quantum Job Scheduling Problem, which has great practical value in the NISQ era. A novel scheduling method is proposed to tackle this problem. With our priority score and noise-aware initial mapping, our method surpasses baselines in time metrics, and guarantees the fairness and fidelity. The contributions of this paper are:
1) We formulate the Quantum Job Scheduling Problem of reducing the latency of (superconducting) quantum processors, fully utilizing the computational power of quantum processors.
2) We propose a novel noise-aware quantum job scheduler to balance time metrics, fidelity, and fairness. The small overhead caused by our method can be neglected.
3) Experimental results on both the noise model and real-world quantum computer show that our approach significantly reduces the QPU time and turnaround time at a low cost of fidelity.
Figure 3. A quantum circuit (left) and its directed acyclic graph (right).
@@ 2. Preliminaries and Related Works
Quantum Computing. The basic unit in QC is a qubit, which is in superposition of basis states |$\ket{0}$ and |$\ket{1}$: |$\ket{\psi}=a\ket{0}+b\ket{1}$, where $|a|^{2}+|b|^{2}=1$. Likewise, a quantum system with $n$ qubits is in superposition of $2^{n}$ basis states. The evolution of quantum states can generate solutions to specific problems, perhaps much faster than classical methods. We refer readers to (Nielsen and Chuang, 2010) for detailed backgrounds. Quantum circuits are employed to implement quantum computation of quantum states. Each quantum circuit consists of quantum gates like X gates, RZ gates, CNOT gates, etc. To obtain the result, we have to repeat executing the circuit many times (shots), because the quantum measurement will cause the collapse of a superposition state to a basis state. A three-qubit quantum circuit is given in Fig. 3 as an example. A quantum circuit can further be converted into a directed acyclic graph (DAG). The topological order of the DAG corresponds with the execution order of quantum gates (from left to right). For example, gate $g_{4}$ cannot be executed until gate $g_{1}$, $g_{2}$ and $g_{3}$ are executed in Fig. 3.
@@
Quantum Processors. The core of a quantum computer is the quantum processor, aka QPU, which serves to execute quantum circuits. We focus on superconducting quantum processors in this paper. The major properties of a QPU are its basis gates, coupling graph and noise condition. Basis gates are the quantum gates supported on the QPU. All the gates in a quantum circuit must be converted to combinations of basis gates during compiling before execution. As shown in Fig. 1, the coupling graph restricts the connectivity of qubits. Two-qubit gates can only be deployed on connected qubits. Besides, the noise of QPUs in the NISQ era results in gate errors, measurement (readout) errors, and decoherence, which will corrupt the quantum state and reduce the fidelity. These errors change over time, so they must be calibrated regularly. Nowadays, many quantum processors are open to public through quantum cloud services. Our submitted quantum jobs will queue up to be executed. In this paper, we conduct experiments on the Qiskit (Aleksandrowicz et al., 2019) noise model of 16-qubit IBM Guadalupe (Fig. 1b), and 66-qubit Xiaohong [^1] As used in our experiments, Xiaohong is a 66-qubit superconducting quantum processor, which can be accessed via public cloud at https://quantumctek-cloud.com/. The used QCIS instruction set can be easily converted from or to the widely used QASM. (Fig. 1a) quantum processor from QuantumCTek (Platform, 2022).
@@
Qubit Mapping. When logical qubits of a quantum circuit are mapped to physical qubits on a QPU, the original two-qubit gates may violate the connectivity constraints as shown in Fig. 2(b). A traditional way to solve this problem is to insert SWAP gates. A SWAP gate is implemented by three CNOT gates (Fig. 1c), incurring extra noise. Hence, the number of them is expected to be minimized. Siraichi et al. formally introduce the aforementioned qubit allocation (mapping) problem (Siraichi et al., 2018), which is proved to be NP-complete. Li et al. propose a bidirectional heuristic search (SABRE) to tackle this problem. When inserting a SWAP gate, they consider its impact on two-qubit gates in both the front layer and extended set, significantly reducing the SWAP overhead (Li et al., 2019). Niu et al. take the error rate and execution time of CNOT gates into consideration, and provide BRIDGE gates as an alternative to SWAP gates (Niu et al., 2020). The BRIDGE gate (Fig. 1d) is composed of four CNOT gates, but its effect equals a single CNOT gate, without changing the mapping. Niu et al. further ameliorate the mapping method by involving the cost of inserted SWAP gates and BRIDGE gates themselves (Niu and Todri-Sanial, 2023). Other methods like Reinforcement Learning (RL) (Huang et al., 2022), Monte Carlo Tree Search (MCTS) (Zhou et al., 2020; Sinha et al., 2022), binary integer programming (Nannicini et al., 2022), and Satisfiability Modulo Theory (SMT) (Murali et al., 2019) have also been studied in qubit mapping. We refer to (Ge et al., 2024) for a more comprehensive survey.
@@
Quantum Multi-Programming (QMP). QMP means running multiple quantum circuits simultaneously on a QPU. This task can be decomposed into two sub-tasks, i.e. qubit partition and qubit mapping. Qubit partition allocates a unique region on the QPU to every parallel quantum circuit. Then, qubit mapping pairs logical qubits with physical qubits on the partition, and inserts SWAP gates to satisfy all two-qubit constraints. Das et al. propose QMP on NISQ devices to improve throughput (Das et al., 2019). They allocate less noisy physical qubits to logical qubits with higher utility. Qucloud (Liu and Dou, 2021) leverages FN community detection algorithm (Newman, 2004) to partition QPUs, and designs an EPST score to estimate the fidelity of allocation. Different quantum circuits can be executed together only when the gap between co-located EPST and separate EPST is less than the threshold. Resch et al. run multiple QAOA (Farhi et al., 2014) circuits in parallel to accelerate the training process (Resch et al., 2021). They greedily expand the partition by breadth-first search (BFS) based on heuristics. All the three methods (Das et al., 2019; Liu and Dou, 2021; Resch et al., 2021) utilize SABRE (Li et al., 2019) to conduct qubit mapping. Niu et al. reorder the quantum circuits according to their CNOT density, and partition QPUs based on the connectivity and error rates of physical qubits (Niu and Todri-Sanial, 2023). These existing methods either disregard the execution order or just focus on QMP in single execution.
@@ 3. Methodology
In this section, we formally introduce the Quantum Job Scheduling Problem (QJSP) to excavate the importance of the execution order when multi-programming massive quantum circuits in the queue of quantum cloud services. Also, a noise-aware quantum job scheduler (NAQJS) is proposed to tackle this problem.
3.1. The Quantum Job Scheduling Problem
@@ 3.1.1. Definition
Suppose the current job queue $\mathcal{Q}$ is comprised of $K$ quantum jobs to be executed, i.e. $\mathcal{Q}=\{\mathcal{J}_{1},\mathcal{J}_{1},\cdots,\mathcal{J}_{K}\}$. Each job $\mathcal{J}_{i}$ can be represented as a tuple $(c_{i},s_{i},t_{i})$, where $c_{i}$, $s_{i}$, and $t_{i}$ denote the quantum circuit, number of measurement shots, and submission time, respectively. Then, new jobs $\mathcal{J}_{K+1},\mathcal{J}_{K+2},\cdots$ will be submitted at time $t_{K+1},t_{K+2},\cdots$. For a quantum computer, besides the execution time $t_{e}$ of circuits on the QPU, other procedures like circuit verification, generation of control signals, and communication will cost extra time $t_{m}$ between execution.
Given the coupling graph $\mathcal{G}$, noise calibration data $\mathcal{N}$, and basis gate set $\mathcal{B}$ of the QPU, we need to execute all the jobs submitted during a time period on the QPU. The objective of QJSP is to minimize the execution latency of jobs, and maintain high fidelity.
@@ 3.1.2. Metrics
The performance assessment of QJSP is divided into two parts: time and fidelity. Also, fairness should be considered.
Time. For users, they mainly care about the time cost from submission to completion of their quantum job, which we name turnaround time. For suppliers, they emphasize on the QPU time of their quantum processors, i.e. total circuit execution time on QPU.
Fidelity. The real fidelity of a quantum state is hard to obtain on a quantum computer, because recovering the complete quantum state from measurements is non-trivial. Methods like classical shadow (Aaronson, 2018; Huang et al., 2020) can mitigate this problem but still incur additional overhead on quantum processors to achieve ideal accuracy. By convention, we use the Probability of Successful Trial (PST), which is defined as the percentage of trials producing the correct result, as our fidelity metric. This metric is widely used in NISQ applications (Linke et al., 2017; Tannu and Qureshi, 2019b; Das et al., 2019; Tannu and Qureshi, 2019a; Liu and Dou, 2021; Niu and Todri-Sanial, 2023) as an alternative to fidelity for its cost-efficiency.
Figure 4. Overview of our noise-aware quantum job scheduler (NAQJS).
@@ 3.2. Proposed Method
Our proposed method is composed of three parts: queue rearranging, qubit partitioning, and qubit mapping. Fig. 4 shows the overview of NAQJS. In each iteration, we sort the current updated queue by our priority score $S_{p}$. Then, we evaluate quantum jobs in the sorted queue one by one, selecting and mapping those jobs whose circuits can find a partition on the remaining coupling graph, until the number of used physical qubits exceed the limit, i.e. $\Sigma_{i=1}^{k}n_{i}\leq\eta\cdot N$, where $n_{i}$ is the number of qubits (i.e. width) of the $i$-th selected jobs, and $k$ is the number of selected jobs. $N$ denotes the number of physical qubits, and $\eta\in(0,1]$ is the allowed maximum usage of physical qubits, which influences the average fidelity because higher usage will incur more noise. Also, $\eta$ can prevent jobs from using extremely noisy qubits. Then, the mapped jobs will be executed in parallel on the QPU. The number of shots is set as the maximum shot number among the mapped jobs to ensure all the shot requirements are satisfied, because those jobs whose shot requirements are unsatisfied will lead to extra execution overhead in following iterations. Since we can retain only the first $s_{i}$ outcomes, this execution pattern will not affect the result. The mapped circuits are executed in an As Late As Possible (ALAP) manner so that circuits with different depth can be measured and completed at the same time to avoid decoherence.
@@ 3.2.1. Queue Rearranging
In each iteration, the job queue will be updated due to new submitted jobs and executed jobs. Akin to the importance of process scheduling for CPUs, the execution order of quantum jobs also counts in QJSP. Therefore, we sort quantum jobs in the updated queue in descending order of priority score. Three properties of a quantum job $\mathcal{J}_{i}$ are considered for its priority score $S_{p}^{(i)}$: the number of qubits $n_{i}$, number of shots $s_{i}$, and submission time $t_{i}$. The priority score is defined as the linear combination:
(1) $S_{p}^{(i)}=-\alpha\cdot S_{n}^{(i)}-\beta\cdot S_{s}^{(i)}-\gamma\cdot S_{t}^{(i)},$
where $\alpha$, $\beta$, and $\gamma$ ($\alpha,\beta,\gamma\geq 0$) denote the width weight, shot weight, and time weight. $S_{n}^{(i)}$, $S_{s}^{(i)}$, and $S_{t}^{(i)}$ are the Min-Max Normalization results of $n_{i}$, $s_{i}$, and $t_{i}$. For example, $S_{n}^{(i)}$ can be calculated as $S_{n}^{(i)}=(n_{i}-n_{min})/(n_{max}-n_{min})$.
The number of qubits $n_{i}$, number of shots $s_{i}$, and submission time $t_{i}$ are deemed as three most important factors for the time and fidelity metric in QJSP. The reason is as follows:
# Qubits. When $n_{i}$ is small, the QPU can accommodate more jobs, which means more jobs are executed in unit time. This is similar to the Shortest Job First (SJF) strategy in process scheduling, which significantly raises the throughput at the beginning, thus improving the average turnaround time.
# Shots. The term $S_{s}^{(i)}$ narrows the distance of $s_{i}$ between neighboring jobs in the queue. Since the shot number is set as the maximum of $s_{i}$ among jobs in one execution, $S_{s}^{(i)}$ can decrease the number of unnecessary shots, which in turn reduces the QPU time. Also, a small $s_{i}$ accelerates the execution of quantum jobs at the beginning, which can benefit the turnaround time.
Submission time. The term $S_{t}^{(i)}$ prioritizes early-submitted jobs, sacrificing the turnaround time and QPU time for fairness, which embodies in the maximum and standard deviation of turnaround time.
It is worth mentioning that the circuit depth is excluded from the calculation of the priority score. For Xiaohong, the execution time of every shot is set as a constant (i.e. 0.2 ms) in reality, no matter how deep the executed circuit is. The constant execution time makes it convenient for the system to operate. Also, this time is long enough for both the execution of the deepest circuit allowed and qubit de-excitation. We have executed circuits of different depth on Xiaohong, and found that the execution time of every shot is about 0.2 ms for all circuits. Similar property is found on IBM quantum cloud empirically. Hence, the circuit depth makes no difference on latency in practical settings.
Alike process scheduling, QJSP also faces the starvation problem (Tanenbaum, 2009) that a quantum job waits infinitely long to run because its priority score is lower than others all the time. Starvation occurs when the number of qubits and shots of a quantum job is extremely large. The term $S_{t}^{(i)}$ mitigates the starvation problem but cannot avoid it. Hence, we adopt an aging strategy, i.e. raising the priority score $S_{p}^{(i)}$ by 1 every $\Delta t$ seconds when job $\mathcal{J}_{i}$ waits in the queue.
Theorem 3.1.
With our aging strategy, all the quantum jobs can be executed in finite time in QJSP.
Proof.
For any job $\mathcal{J}_{i}$ after the Min-Max Normalization, $S_{n}^{(i)}$, $S_{s}^{(i)}$ and $S_{t}^{(i)}$ range from 0 to 1. Then, the priority score $S_{p}^{(i)}\in\left[-\alpha-\beta-\gamma,0\right]$. When the time reaches $t^{\prime}\coloneqq t_{i}+(\alpha+\beta+\gamma+1)\Delta t$, its priority score satisfies $S_{p}^{(i)}\in\left[1,\alpha+\beta+\gamma+1\right]$, larger than that of any job submitted after $t^{\prime}$. Hence, any job submitted after $t^{\prime}$ will be executed later than $\mathcal{J}_{i}$. Since the number of jobs submitted before $t^{\prime}$ is finite, $\mathcal{J}_{i}$ will be executed in finite time. ∎
Input: DAG of the Circuit $G$, Number of Qubits $n$
Output: Circuit Time $t_{c}$
1 Initialize $t[i]=0$, for $i=0,1\cdots n-1$;
2 Initialize an empty queue $Q$;
3 for gate $g$ in $G$ do
4       if $g.in\_degree==0$ then
5             $Q.push(g)$;
6
7       end if
8
9 end for
10while $Q$ is not empty do
11       $g=Q.top()$;
12       if $g$ is one-qubit gate then
13             $q_{1}=g.qubits$;
14             $t[q_{1}]=t[q_{1}]+g.duration$;
15
16       end if
17      if $g$ is two-qubit gate then
18             $q_{1},q_{2}=g.qubits$;
19             $t[q_{1}]=t[q_{2}]=max(t[q_{1}],t[q_{2}])+g.duration$;
20
21       end if
22      for gate $g^{\prime}$ in $g.successors$ do
23             $g^{\prime}.in\_degree=g^{\prime}.in\_degree-1$;
24             if $g^{\prime}.in\_degree==0$ then
25                   $Q.push(g^{\prime})$;
26
27             end if
28
29       end for
30      $Q.pop()$;
31
32 end while
[…]
34
Algorithm 1 Calculation of circuit time